Phyllis invested 39000 dollars, a portion earning a simple interest rate of 4 percent per year and the rest earning a rate of 6 percent per year. After one year the total interest earned on these investments was 1980 dollars. How much money did she invest at each rate?

Answer by geetha_rama(94)

Let X be the amount invested at 4% per year
and 39000 - X be the amount at 6% per year.
Total interest earned = 1980
Using simple interest formula,
(X*4)/100 + (39000 -X)*6/100 = 1980
Diving by 2,
(X*2)/100 + (39000-X) * 3/100 = 990
2X + 117000 - 3X = 99000
117000 - 99000 = X
=>X = 18000
39000-X = 21000
Amount invested at 4% per year = $18000
Amount invested at 6% per year = $21000